Discusses clearly macroeconomic equilibrium condition by using an appropriate graphs and both aggregate demand & aggregate supply equation.

1
Expert's answer
2022-03-21T13:51:16-0400

Solution.

Equilibrium refers to when price adjust upwards or downward to achieve balance

I)increase in demand



SS is the supply curve and D1D1 the initial demand curve shifts to the right, to position D2D2. P1 is the initial equilibrium price and q1 the initial equilibrium quantity. When demand increases to D2D2, then at price P1, the quantity demanded increases from q1 to qd. But the quantity supplied at that price is still q1. This leads to excess demand over supply (qd – q1). This causes prices to rise to a new equilibrium level P2 and the quantity supplied to rise to a new equilibrium level, q2. ii). Decrease in Demand

At the initial equilibrium price P1, quantity demanded falls from q1 to qd. But the quantity supplied is still q1 at this price. Hence, this creates excess of supply over demand, and this causes price to fall to a new equilibrium level P2 and quantity to fall to a new equilibrium level

iii).increase in supply


DD is the demand curve and S1S1 the initial supply curve. If supply increases, the supply curve shifts to the right to position S2S2. At the initial equilibrium price P1, quantity supplied increase from q1 to q2. This creates a glut in the market and this causes the price to the new P21 and the quantity increases to a new equilibrium level q2.

iv).fall in supply

When the supply falls, the supply curve shifts to the left to position S1S1. At the initial equilibrium price P1, quantity supplied falls from q1 to q21 but the quantity demanded is still q1. This creates excess of demand over supply which causes price to rise to a new equilibrium level P21 and quantity to fall to a new equilibrium level q21 and quantity to fall to a new equilibrium level q2.


Aggregate demand refers to the total planned spending in the economy as a whole in a given period. It is made up of consumption demand by individuals, planned investment demand, government demand and demand by foreigners of the nationals output.

The equation is given as AD = C + I + G +(X-M)

Where AD =aggregate demand, C is consumption, I is investment, G is the government spending, X is the total exports and M is total imports.


Aggregate supply is the total foods firms are able to supply in an economy in a given period.

The equation is Y = Y* + "\alpha" (P - P"\epsilon" )

where Y is the economy production, Y* is the natural level of production, "\alpha" is the coefficient ,P is the price level, P"\epsilon" is the expected price level from consumers.